OpenStudy (anonymous):

square root of a - 4 / 3 square root of a - square root of b

OpenStudy (anonymous):

can u give it in equation form

OpenStudy (anonymous):

\[\sqrt{a}-4\div3\sqrt{a}-\sqrt{b}\]

OpenStudy (anonymous):

\[(3a-12\sqrt{a}-4\sqrt{b}+\sqrt{ab})/(9a-b)\]